OpenStudy (anonymous):

2. x^4+4 is equal to (a) (x^2+2)(x^2-2) (b) (x^2+2x)(x-2) (c) (x^2+2x+2)(x^2-2x+2) (d) none of these 2. x^4+4 is equal to (a) (x^2+2)(x^2-2) (b) (x^2+2x)(x-2) (c) (x^2+2x+2)(x^2-2x+2) (d) none of these @Mathematics

OpenStudy (anonymous):

(d)

OpenStudy (anonymous):

sorry, (C)

OpenStudy (anonymous):

Can you explain how?

OpenStudy (anonymous):

I think it can be rewritten as (x^2)^2+2^2

OpenStudy (anonymous):

yeah, add 4x^2 and again minus 4x^2 u will get the expression :)

OpenStudy (anonymous):

Did not understand.

OpenStudy (anonymous):

x^4+4=x^4+4x^2+4-4x^2=(x^2+2)^2-(2x)^2=(x^2+2x+2)(x^2-2x+2)

OpenStudy (anonymous):

How (x^2+2)^2-(2x)^2=(x^2+2x+2)(x^2-2x+2) ?

OpenStudy (anonymous):

because, x^2-y^2=(x+y)(x-y)

OpenStudy (anonymous):

Thanks Arnab

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

What does np mean?

OpenStudy (anonymous):

no problem :)